Selected financial information for Company B is shown below for the fiscal years 2018 and 2019.

2018 2019
Revenue $152,256 $155,427
Gross Profit $12,020 $20,502
Operating Income $(30,363) $5,231
Net Income $6,188 $5,346
Total Assets $149,422 $166,344
Total Liabilities $112,422 $123,170

What are the Leverage and Return on Equity values for 2019? Use the average of beginning and ending balances for balance sheet items.

A. 0.98 and 51%, respectively

B. 1.34 and 13%, respectively

C. 1.34 and 51%, respectively

D. 3.94 and 13%, respectively
Answer #1

Average net income= ($6188+$5346) /2 = $5767

Average total assets= ($149422+$166344) /2 = $157883

Begining Shareholders equity= Begining total assets - Begining total liabilities = $149422 - $112422 = $37000

closing Shareholders equity= closing total assets - Closing total liabilities = $166344 - $123170 = $43174

Average shareholders equity = ($37000 + $43174) /2 = $40087

Leverage = Average Total Asset / Average Equity = $157883 / $40087 = 3.9385 i.e. 3.94

Return on Equity = Net income of 2019 / Average equity = ($5346 / $40087) *100 = 13.33% i.e. 13%

Therefore, we take the answer of D i.e. Leverage and ROE are 3.94 and 13%, respectively.